### v2

# A three-state coupled Markov switching model for COVID-19 outbreaks across Quebec based on hospital admissions

Abstract
Recurrent COVID-19 outbreaks have placed immense strain on the hospital system in Quebec. We develop a Bayesian three-state coupled Markov switching model to analyze COVID-19 outbreaks across Quebec based on admissions in the 30 largest hospitals. Within each catchment area, we assume the existence of three states for the disease: absence, a new state meant to account for many zeroes in some of the smaller areas, endemic and outbreak. Then we assume the disease switches between the three states in each area through a series of coupled nonhomogeneous hidden Markov chains. Unlike previous approaches, the transition probabilities may depend on covariates and the occurrence of outbreaks in neighboring areas, to account for geographical outbreak spread. Additionally, to prevent rapid switching between endemic and outbreak periods we introduce clone states into the model which enforce minimum endemic and outbreak durations. We make some interesting findings, such as that mobility in retail and recreation venues had a positive association with the development and persistence of new COVID-19 outbreaks in Quebec. Based on model comparison our contributions show promise in improving state estimation retrospectively and in real-time, especially when there are smaller areas and highly spatially synchronized outbreaks. Furthermore, our approach offers new and interesting epidemiological interpretations, such as being able to estimate the effect of covariates on disease extinction.
Key words : Bayesian inference; COVID-19; Hidden Markov model; Outbreak detection; Outbreak forecasting; Zero-inflation.
1 Introduction
Quebec has been the epicenter of the COVID-19 epidemic in Canada with 15,389 deaths and 52,788 hospitalizations through May 2022 (INSPQ, 2022). As a result, immense strain has been placed on Quebec’s hospital system. Due to the high demand for hospital beds, many elective surgeries have been delayed or canceled which can adversely affect outcomes in patients with conditions other than COVID-19 (Shingler and Hendry, 2022). Additionally, there have been concerning shortages of essential medical supplies, drugs and staff (Laframboise, 2020; Legault and Blais, 2020).
In this paper we analyze weekly COVID-19 hospital admissions in the 30 largest hospitals in Quebec, a map is given in Figure 1 along with the time series for each hospital. Figure 2(a) shows the weekly COVID-19 hospitalizations for Pavillon Sainte-Marie in Trois–Rivières, an average-sized hospital among the 30. As illustrated in Figure 2(a), and this can also be observed less clearly in Figure 1 across all hospitals, the hospital admissions can be characterized by a series of outbreak periods separated by quiescent endemic periods with low levels of hospitalizations. The outbreak periods represent COVID-19 outbreaks in the surrounding communities serviced by the hospital (catchment areas) as we do not include hospital-acquired infections. Clearly, most hospitalizations occur during the outbreak periods. However, this does not reflect the true burden of the outbreaks as the rapid rise in hospitalizations at the beginning of an outbreak can be difficult to adjust to, leading to shortages and a lack of beds (Shingler and Hendry, 2022). To gain a better understanding of how the outbreaks develop, a substantive aim of our analysis is to quantify how certain factors, such as mobility and the introduction of new variants (marked in Figure 2(a), see Section 5.1 for more details), are associated with the emergence and persistence of the outbreaks. Additionally, we want to detect and forecast the outbreaks so that the hospitals can better plan the allocation of their medical resources.
Many methods have been proposed for modeling infectious disease counts in space and time, a subject that has gained increased attention in recent years partly due to the COVID-19 pandemic. Among the most popular are epidemiological compartmental models, such as susceptible-infectious-recovered (SIR) models (Keeling and Rohani, 2007; Bjørnstad, 2023). These approaches attempt to model individuals within the population moving through different disease compartments such as susceptible, infectious, recovered and quarantined (Crawford et al., 2022). Multivariate autoregressive count time series models, where the current expected counts are modeled as some function of past counts, are also popular (Ssentongo et al., 2021; Bracher and Held, 2022) and can be motivated from discrete-time SIR models (Bauer and Wakefield, 2018; Mizumoto and Chowell, 2020). When fit with covariates, both of these approaches are often used to investigate associations between certain factors and some measure of disease transmission, e.g., the average number of new cases per each recent previous case (Bauer and Wakefield, 2018; Flaxman et al., 2020; Ssentongo et al., 2021). However, they usually do not clearly distinguish between outbreak and non-outbreak periods. As mentioned above, we are interested in studying the transition between calm endemic periods and outbreak periods, and, therefore, we require a precise model definition of those transitions. Also, these methods are not well suited for outbreak detection and forecasting. While a compartmental model could forecast a likely increase in some epidemiological indicator, it cannot distinguish between an increase due to random endemic variation and an actual outbreak developing, which is an important problem in outbreak detection (Unkel et al., 2012). Several other methods that are more focused on producing predictions, such as spline-based (Bauer et al., 2016) or machine learning (Rahimi et al., 2021) methods, are also popular, however, they are often not appropriate for learning about the dynamics of the disease.
In this paper we take a Markov switching approach to modeling the occurrence of the outbreak periods. As we explain below, this approach allows us to study, detect and forecast transitions between endemic and outbreak periods. Markov switching models assume a time series can be described by several submodels, usually called states or regimes, where switching between submodels is governed by a hidden first-order Markov chain (Hamilton, 1989). Markov switching models that switch between endemic and outbreak states have a long history (Souza, 1982; Amorós et al., 2020). In this framework, the transitions between the endemic and outbreak states are usually modeled as a change in the parameters of an autoregressive process, i.e., a change in transmission intensity (Lu et al., 2010). These transitions are assumed unobservable and must be inferred probabilistically from the time series, which is convenient as we usually do not observe exactly when the outbreaks start and end. The probabilities that govern the transitions between the endemic and outbreak states can depend on covariates (Diebold et al., 1994). This allows for investigating the association of various factors with the probabilities of outbreak emergence and persistence and can aid in forecasting the outbreaks (Nunes et al., 2013). Additionally, Bayesian methods can be used to compute the posterior probability an outbreak is currently happening or will happen soon for the purpose of outbreak detection (Martínez-Beneito et al., 2008) and forecasting (Nunes et al., 2013), respectively. Outbreak/endemic Markov switching models have become increasingly popular in epidemiology, especially for outbreak detection (Unkel et al., 2012), with recent applications to influenza (Lytras et al., 2019), cutaneous leishmaniasis (Rahmanian et al., 2021) and salmonella (Zacher and Czogiel, 2022).
Despite their growing popularity, outbreak/endemic Markov switching models have not focused on the analysis of outbreaks in small areas, especially with many zeroes. Figure 2(b) shows weekly COVID-19 hospitalizations for the Hôpital Anna–Laberge in Châteauguay, a relatively small hospital. As illustrated in Figure 2(b), it may be more appropriate in smaller areas to describe switching between three periods: absence, endemic and outbreak. This has a strong epidemiological justification as it is well known that many infectious diseases frequently go extinct in small communities (Bartlett, 1957; Keeling and Rohani, 2007). Ignoring long periods of disease absence would end up likely assigning too many zeroes to the endemic state, biasing its mean towards zero. This is undesirable as it could lead to the outbreak state being too dominant at medium counts, leading to false alarms (Rath et al., 2003). Another advantage of considering an absence state is that we can model the probabilities of disease extinction and reemergence, which are epidemiologically interesting (Douwes-Schultz and Schmidt, 2022).
There has also not been much work done on spatio-temporal outbreak/endemic Markov switching models, where the focus is on the analysis of outbreaks in several different but connected areas across time. This is a focus of our analysis as many important decisions are made at the individual hospital level and clearly, there is some, though not perfect, synchronization of outbreaks between hospital catchment areas, compare Figures 2(a) and 2(b) for example. Amorós et al. (2020) fit a spatio-temporal outbreak/endemic Markov switching model, but they only borrowed spatial strength in the observation component of the model. Individuals will mix between areas, causing outbreaks to spread geographically (Grenfell et al., 2001). Therefore, an outbreak should be more likely to emerge or persist in an area if there are outbreaks likely occurring in neighboring areas. One way to achieve this is to use a coupled Markov switching model (Pohle et al., 2021) where the states of neighbors are directly entered into the transition probabilities (Douwes-Schultz and Schmidt, 2022; Touloupou et al., 2020). For outbreak detection and forecasting, considering evidence of outbreaks in neighboring areas could provide an early warning at the very beginning of an outbreak when there is still uncertain evidence within the area. Heaton et al. (2012) did let the probabilities of outbreak emergence depend on outbreaks in neighboring areas, however, they did not allow the transition probabilities to depend on covariates, and they used an absorbing state model, meaning it is difficult to apply their model to time series that contain multiple outbreaks.
In our framework, we assume the disease switches between three states in each area: absence, endemic and outbreak. The epidemiological count in the absence state is assumed to always be 0, while it follows two autoregressive negative binomial processes in the outbreak and endemic states, distinguished by a lower level of transmission during the endemic period. Switching between the three states is governed by a first-order Markov chain where we focus on modeling the following four transition probabilities: absence to endemic (disease emergence), endemic to absence (disease extinction), endemic to outbreak (outbreak emergence) and outbreak to outbreak (outbreak persistence). Each transition probability can depend on covariates as well as a weighted sum of outbreak occurrence in neighboring areas. This allows us to investigate associations with important epidemiological transitions while also incorporating outbreak spread between areas. Nonhomogeneous Markov switching models can be sensitive to overfitting as the transition matrix may be non-persistent at some levels of the covariates, potentially leading to rapid switching between the states. Clearly, it is not realistic to rapidly switch between outbreak and endemic states, so we additionally introduce clone states (Kaufmann, 2018) into the model to enforce a minimum endemic and outbreak duration.
This paper is structured as follows. In Section 2 we introduce our proposed model, a three-state coupled Markov switching model. In Section 3 we describe Bayesian inference using data augmentation and make use of the individual forward filtering backward sampling algorithm of Touloupou et al. (2020). In Section 4 we evaluate outbreak detection and forecasting results from the model on simulated data where the exact start and end times of the outbreaks are known. In Section 5 we apply the model to COVID-19 outbreaks across Quebec based on admissions in the 30 largest hospitals. We close with a general discussion in Section 6.
2 A Three-state Coupled Markov Switching Model
Let be an epidemiological count indicator, e.g., counts of hospitalizations, associated with area across time periods. Let be an indicator for the epidemiological state of the disease, where if the disease is absent in area during time , if the disease is in an endemic state and if the disease is in an outbreak state.
We assume that the epidemiological count indicator is always 0 when the disease is absent and follows two distinct autoregressive negative binomial processes in the endemic and outbreak states, that is,
where and are the means in the endemic and outbreak states respectively and and are the overdispersion parameters, so that, for example, . A zero count could be produced by all three states while a positive count can be produced by either the endemic or outbreak states, therefore, none of the states are observable and is a latent variable. Intuitively, we often do not know if an outbreak is occurring and there could be zeroes during an endemic period, or early outbreak period, due to a failure to detect the disease or a lack of severe cases for hospitalizations and deaths. As justification for the negative binomial distribution in (1), over a Poisson distribution, we found there was a high amount of overdispersion in the Quebec hospitalizations, see Table 3.
As we are modeling an infectious disease we would expect the previous count to affect the current expected count when the disease is present (Bauer and Wakefield, 2018). Therefore, we use log-linear autoregressive models (Liboschik et al., 2017) for and ,
where and are random intercepts meant to account for between area differences and is a vector of space-time covariates that may affect transmission of the disease within the endemic and outbreak periods. We allow the covariate effects, and , to be different as outbreaks can lead to behavioral changes in hosts (Verelst et al., 2016). We also considered spatially correlated random effects (Amorós et al., 2020) for the area specific intercepts in (2). However, we found for our motivating example that, possibly due to the amount of uncertainty about the underlying disease states, the spatial association of the intercepts could not be estimated precisely enough.
To model the switching between absence, endemic and outbreak periods we assume that follows a three-state nonhomogeneous Markov chain within each area. In order to incorporate outbreak spread between areas we condition the transition matrix on , the vector of all state indicators excluding area at time . We propose the following conditional transition matrix for the Markov chain, for ,
where for and we have the following epidemiological interpretations of the transition probabilities,
From (3), we assume it is not possible to move from absence to outbreak and vice versa in a single time step. For a time step of one week, like in our application, this is reasonable. However, for longer time steps this assumption should be examined.
We assume each transition probability in (3) can depend on a -dimensional vector of space-time covariates as well as a weighted sum of outbreak occurrence in neighboring areas during the previous time period, to model outbreak spread between areas. Starting with the probabilities of disease emergence, , and outbreak persistence, , we let
for , where is an indicator function, is the set of all neighboring areas of area and is a fixed weight meant to reflect a priori knowledge of the level of influence area has on area . For example, one could use distance-based weights or weights based on trade between regions, see Schrödle et al. (2012). It is also possible to estimate connectivity in a coupled Markov switching model (Douwes-Schultz and Schmidt, 2022). However, we do not consider this here due to the complexity and number of transitions in our model, the model used in Douwes-Schultz and Schmidt (2022) only had one general presence state and one absence state. For the probabilities of disease extinction, , and outbreak emergence, , a multinomial logistic regression is needed so that the second row of (3) sums to 1, that is,
for . Therefore, and in (5) represent the effects of the covariates on the relative odds of transitioning to the absence and outbreak states respectively compared to remaining in the endemic state.
Covariates will likely affect the transition probabilities in most cases. For example, dengue outbreaks are unlikely to occur during the winter when mosquito activity is low (Descloux et al., 2012), the introduction of new variants, such as Omicron, has likely played an important role in the development of new COVID-19 waves (Maslo et al., 2022) and disease extinction is less likely in smaller communities (Bartlett, 1957). As for the inclusion of neighboring outbreak indicators in (4)-(5), individuals will mix with those in other areas which can cause outbreaks to spread geographically, a phenomenon known as traveling waves (Grenfell et al., 2001). Therefore, an outbreak in a neighboring area should affect the transition probabilities due to either direct spread or because it indicates spread from a common source, such as a large city.
2.1 Identifiability constraints
Constraints are often placed on the parameters of a Bayesian mixture model to remove multimodality in the posterior distribution (Frühwirth-Schnatter, 2006), which often occurs genuinely and due to the invariance of the likelihood function to permutations in the state labeling of the parameters (Jasra et al., 2005). The constraints are commonly chosen based on knowledge of the process being modeled (Martínez-Beneito et al., 2008; Stoner and Economou, 2020). To help motivate constraints in our case, note that another way to express a log-linear model in (2) is as, for example, , which takes the form of a transmission rate multiplied by the previous counts. Disease transmission should always increase when moving from the endemic state to the outbreak state, especially if the model is to be used for issuing alarms for surveillance. Additionally, the autoregressive coefficients, and , control the speed the disease moves through the local population (Wakefield et al., 2019), which should be higher during the outbreak periods. Therefore, we assume the following constraints,
meaning we are assuming at least a one percent increase in transmission when moving from the endemic state to the outbreak state (in Section 5.2 we show the results are not very sensitive to reasonably changing this one percent value). Initially, we had considered a simpler constraint on just the intercepts and the autoregressive coefficients, however, we found in a simulation study, in Section 2 of the supplementary materials (SM), that the simpler constraint does not ensure consistent convergence of our Markov chain Monte Carlo (MCMC) algorithm. Constraining the entirety of the transmission rates, as in (6), greatly improved the convergence rate, did not introduce any significant bias into the inferential procedure and did not add a substantial amount of time to the model fitting.
2.2 Prior specification
For the count part of the model, we specified wide independent normal and gamma priors for most lower-level elements of . We used priors for and to meet the stability conditions of Liboschik et al. (2017) which ensures the counts are not expected to grow without bound. For and we used and priors respectively. An upper limit of 10 was chosen for as it is important for the endemic state distribution to have a long right tail to prevent frequent false alarms during outbreak detection (Rath et al., 2003). To impose the constraints in (6) we truncated the prior distribution of (Kaufmann, 2018). Some shrinkage to the null is generally recommended for logistic and multinomial logistic regression parameters to avoid separation issues and reduce bias away from 0 (Bull et al., 2002). Therefore, following Gelman et al. (2008), we used Cauchy priors with scale for the effects of covariate on the transition probabilities in (4)-(5). For the effects of neighboring outbreaks on the transition probabilities, i.e., for in (4)-(5), we used more aggressive shrinkage and assigned priors. This states that with probability .95 a priori we believe that an outbreak occurring in a single neighboring area should not more than double or less than halve the odds, or relative odds, of any epidemiological transition, e.g., should not more double the odds of an outbreak emerging relative to remaining in the endemic state (Wakefield, 2013). Our reasoning is that we do not want a single area to be given too much influence as we want the model to consider evidence of outbreaks across multiple neighboring areas. Additionally, while borrowing spatial strength can be important, we do not want it to overpower within-area information too strongly.
2.3 Clone states
As the transition probabilities in (3) depend both on covariates and latent neighboring states, the transition matrix may fluctuate between persistence and non-persistence. This could lead to time periods where there is rapid switching between the outbreak and endemic states, which is not realistic. Heaton et al. (2012) dealt with this by using an absorbing state model where the probability of outbreak persistence was fixed at 1, but that does not allow the analysis of multiple outbreaks or the study of outbreak persistence. Our solution is to introduce clone states into the model with determined transitions to enforce a minimum endemic and outbreak state duration, which is common in econometrics (Kaufmann, 2018). We can introduce a new latent state indicator such that if , if and if , and with the following conditional transition matrix, for ,
where for . The new transition matrix (7) will prevent rapid switching between endemic and outbreak periods by imposing a minimum endemic duration of 2 weeks and a minimum outbreak duration of 4 weeks for our motivating example. A COVID-19 outbreak that lasts less than 4 weeks in Quebec is likely either a false alarm or too small to be concerning. Also, there should be at least some time between outbreaks for the susceptible population to replenish (Keeling and Rohani, 2007). Clearly, the idea of (7) could be used to impose any arbitrary minimum state durations, and so we assume (7) throughout the rest of the paper w.l.o.g..
We will refer to the model defined by (1)-(5), with (3) replaced by (7), as the coupled Markov switching negative binomial model with 1 absence state, 2 endemic states and 4 outbreak states, i.e., the CMSNB(1,2,4) model. To finish model specification we also need to specify an initial state distribution for the Markov chain in each area, i.e., for , which we assume does not depend on any unknown parameters. Note that moving from (3) to (7) does not add any new parameters to the model, however, the restrictive transition matrix does slightly complicate the inferential procedure, which will now be discussed.
3 Inferential Procedure
Let be the vector of all state indicators, let be the vector of all counts, let be the vector of all model parameters in the count part of the model, i.e., parameters in (1)-(2), let be the vector of all model parameters in the Markov chain part of the model, i.e., parameters in (4)-(5), and, finally, let be the vector of all model parameters. Then the likelihood of given and is given by,
Recall from the previous section that is not observed. It is not possible to marginalize out from (8) as doing so would require matrix multiplication with a matrix (Douwes-Schultz and Schmidt, 2022). Additionally, we want to make inferences about for the purpose of outbreak detection and forecasting as well as historical retrospection. Therefore, we estimate along with by sampling both from their joint posterior distribution which, from Bayes’ theorem, is proportional to,
where is the prior distribution of .
As the joint posterior (9) is not available in closed form, we resort to MCMC methods, in particular, we used a hybrid Gibbs sampling algorithm with some steps of the Metropolis–Hastings algorithm to sample from it. We sampled all elements of without conjugate priors individually, using an adaptive random walk Metropolis step (Shaby and Wells, 2010). It is easy to implement the sampling of each element of one-at-a-time from (Douwes-Schultz and Schmidt, 2022). However, we found that one-at-a-time sampling mixed so slowly that it is not usable with our model. To see this, consider the following hypothetical state sequence for , . Note it is not possible to sample any new single element of this sequence and in general we found that one-at-time sampling gets stuck in small regions of the parameter space. To avoid issues like this with one-at-a-time sampling when fitting Markov switching models, Chib (1996) proposed to sample all of jointly from . However, this is not possible with our model as it would involve matrix multiplication with matrix (Douwes-Schultz and Schmidt, 2022). As an alternative, we can block sample with each block containing all the state indicators in a single location (Touloupou et al., 2020). Let denote the vector of all state indicators in area and let be with removed. Then we can sample all of jointly from its full conditional distribution, which is given by,
using an individual forward filter backward sampling (iFFBS) algorithm (Touloupou et al., 2020). More details are given in SM Section 1. It is also possible to block sample in multi-location blocks (Douwes-Schultz and Schmidt, 2022) but we do not consider this here as it does not scale well with large transition matrices.
Our hybrid Gibbs sampler was implemented using the R package Nimble (de Valpine et al., 2017). Nimble comes with built-in Metropolis–Hastings and categorical (equivalent to one-at-a-time sampling) samplers. The iFFBS samplers were implemented using Nimble’s custom sampler feature. All Nimble R code, including for the custom iFFBS samplers, are provided on GitHub (https://github.com/Dirk-Douwes-Schultz/CMSNB124_code). Nimble was chosen as it is extremely fast (C++ compiled) and only requires the coding of new samplers. In the SM Section 2, we provide a simulation study, which shows that our proposed Gibbs sampler can recover the true parameters of a CMSNB(1,2,4) model that is specified like in our motivating example in Section 5.
3.1 Outbreak detection, forecasting and historical retrospection
Once a sample from the joint posterior (9) has been obtained, the posterior probability that the disease is currently in epidemiological state , for (absence), (endemic) and (outbreak), in area can be approximated as , where the superscript denotes a draw from the posterior distribution of the variable, is the size of the burn-in sample and is the total MCMC sample size. The posterior probability represents the probability that an outbreak is currently happening in area given all observed data and can be used for the purpose of outbreak detection (Martínez-Beneito et al., 2008). We recommend using along with other, external, pieces of information to help determine if an outbreak is likely occurring. While has no closed form, we can condition on the parameters and previous states to gain some intuition into how the model performs outbreak detection,
From (11), the model generally weighs two factors for outbreak detection, the relative likelihood that the outbreak state generated the observed count and the probability of entering the outbreak state at the current time , which recall may depend on outbreaks in neighboring areas and covariates, see (4)-(5) and (7).
As for forecasting, we used a simulation procedure to draw realizations from the posterior predictive distributions (Frühwirth-Schnatter, 2006). Algorithm 1 in the SM will obtain realizations from the posterior predictive distribution of the counts, , and the epidemiological state of the disease, , for , and . Then, the posterior probability that the disease will be in state , time steps from now, in area , can be approximated as .
Finally, it is also important to examine the posterior probability that the disease was in state during past time periods for , and (absence), (endemic) and (outbreak) for several reasons. Firstly, it may be of historical interest to investigate the epidemiological history of the disease in various areas. Also, the estimation of depends heavily on the estimates of the past states as indicated by the joint likelihood function (8). If the classification of past epidemiological periods is not sensible, this would cast doubt on the estimates of and may point to model misspecification.
3.2 Model comparison
To compare competing models, we use the widely applicable information criterion (WAIC) (Gelman et al., 2014). For state-space models, the WAIC is more accurate when the latent states are marginalized (Auger-Méthé et al., 2021). For models without neighboring outbreak indicators in the transition probabilities, such as the Non-coupled Model in Section 5, we can use the forward filter (Frühwirth-Schnatter, 2006) to calculate the marginal density , where , and use it to calculate the WAIC. However, for models with neighboring outbreak indicators in the transition probabilities it is not computationally possible to completely marginalize (Douwes-Schultz and Schmidt, 2022) and, therefore, we use the partially marginalized density instead. See SM Section 3.1 for more details. The model with the lowest WAIC is considered to have the best fit and as a rule of thumb, a difference of 5 or more in the WAIC is considered significant. A simulation study we conducted, in SM Section 3.2, suggests that when there is no significant difference in the WAIC to prefer the less complex model and when there is a significant difference the WAIC almost always chooses the correct model, between a spatial and non-spatial Markov switching model.
4 Simulation Study
In Sections 5.3 and 5.4 below we attempt to evaluate the retrospective and real-time state estimates of a CMSNB(1,2,4) model (see Section 3.1) fitted to the Quebec hospitalizations. In this case, it is difficult to exactly quantify the accuracy of the state estimates as the true underlying states are not known, and we can only evaluate the estimates visually. One solution is to design a simulation study where the exact start and end times of the outbreaks are pre-determined, although this does have the disadvantage of using simulated and not real data (Buckeridge, 2007).
We designed a simulation study with 30 areas divided into 5 clusters of 6 areas each. Each area alternated between a 15-week endemic period and a 15-week outbreak period for a total of 4 of each period, 120 weeks total. For every area in a cluster, the start of each outbreak was randomized to occur within the first 4 weeks of the corresponding outbreak period. Therefore, we did not explicitly model outbreak spread within a cluster, we just assumed the outbreaks occurred around the same time. We also assumed a 40% chance, taken roughly from our results in Section 5.2, that each endemic period contains a 7-week absence period inserted into the middle. The counts were simulated from a negative binomial distribution with overdispersion 10 and conditional means (endemic), (outbreak) and (absence). The covariate was taken from our real data example to produce some between-area heterogeneity in the transmission. These values are somewhat similar to those from Section 5.2 and produced realistic looking simulations, SM Figure 4 shows some of the simulated time series. We only generated one simulation as each model must be fit 20 times to produce the needed state estimates, see below.
To fit to the simulated data, we considered a CMSNB(1,2,4) model with the count part correctly specified, including no random intercepts, and included only in the relative odds of outbreak emergence, where contained all areas in the same cluster as area . We will call this the Spatial Model, and for comparison purposes, we also considered a homogeneous non-spatial model without in the transition probability. We focused on outbreak emergence in the simulation study, as it is the most important transition to capture in a surveillance setting. We evaluated three outbreak state estimates from the models: (1) retrospective probabilities, from fitting the models to the full simulated data set (2) real-time detection probabilities, from fitting the models up to time for (20 separate fits) and (3) real-time one week ahead forecasts, from fitting the models up to time for . We only evaluated the real-time state estimates on the last outbreak in each area to ensure stable model fitting. We evaluated the outbreak state estimates using the area under the ROC curve (AUC), sensitivity, specificity and timeliness (Buckeridge, 2007), the last three calculated with a 50% threshold. Timeliness for a single outbreak is simply the number of weeks into the outbreak, starting at one, when the outbreak state estimate first rises above 50%, and the overall timeliness is the average timeliness across all outbreaks.
The results are summarized in Table 4 of the SM. The Spatial Model was able to recover the true underlying outbreak states well retrospectively, AUC of .995, and during outbreak detection, AUC of .983, and one week ahead forecasting, AUC of .965. The Spatial Model improved, over the Non-spatial Model, every criterion for each type of outbreak state estimate, especially for real-time outbreak detection and forecasting. For instance, the Spatial Model improved the sensitivity of the real-time outbreak detection probabilities by 6% and reduced the timeliness of the one-week ahead outbreak forecasts by one week. Additionally, the Spatial Model showed a fair tradeoff between outbreak detection and forecasting, with the outbreak forecasts having two percent lower sensitivity but an improvement in timeliness of .43 weeks. In contrast, the forecasts from the Non-spatial Model reduced the sensitivity by 6.7% and only improved the timeliness by .06 weeks. This suggests outbreak forecasting should only be attempted by nonhomogeneous models. These points are further illustrated in Figure 3 which compares the real-time outbreak detection probabilities and the one-week ahead outbreak forecasts in one of the areas from the simulation study.
5 Application to COVID-19 Outbreaks Across Quebec
5.1 Model specification and fitting
We fitted a CMSNB(1,2,4) model to weekly COVID-19 hospitalizations across the 30 largest hospitals in Quebec (), in terms of overall COVID-19 admissions, between March 16th 2020 and May 9th 2022 (). Hospitalizations typically lag infections by 1–2 weeks, (Ward and Johnsen, 2021) and there are a few days of reporting delay. We did not include hospital-acquired infections so that the outbreak periods represent outbreaks in the catchment areas of the hospitals. A hospital’s catchment area is typically difficult to exactly determine and can stretch disjointly across a broad geographical region (Gilmour, 2010). By examining the locations of a sample of patients from each hospital, we found that the catchment areas are mostly contained within, and sometimes adjacent to, the same county where the hospital is located. We assumed that the number of beds in a hospital is a reasonable proxy for the population size of the catchment area.
Increased mobility has been a major concern of the Quebec government during the epidemic due to a fear that it will lead to increased COVID-19 transmission, potentially overwhelming hospitals (Wilton, 2020). In general, epidemiologists have long theorized that mobility plays an important role in the development of recurring outbreaks (Soper, 1929). While quantifying mobility is challenging, Google mobility metrics (Google LLC, 2022) tend to give a reasonable approximation to more accurate, but only privately available, mobility measures based on close contact rates (Crawford et al., 2022). Google LLC (2022) provides daily metrics that measure mobility as a percent change in the number of visits to certain venues, based on personal electronic devices, from days of the same weekday in January 2020. We used the retail+recreation mobility metric which measures changes in visits to places like restaurants, cafes, shopping centers, museums, libraries, and movie theaters as such venues have been heavily targeted by the Quebec government (Gouvernement du Québec, 2022). In Quebec, Google mobility metrics are available at the county level, although the metrics are missing for around 10 percent of counties. Another major concern has been the introduction of new COVID-19 variants, especially the Alpha and Omicron variants (Olivier, 2021; Stevenson, 2021). New variants of COVID-19 can be more contagious and more resistant to existing vaccines compared to previous variants, although they can also be less deadly (Chenchula et al., 2022).
Based on the above discussion, we considered as covariates possibly associated with transmission during the endemic and outbreak periods in Equation (2) and disease emergence, disease extinction, outbreak emergence and outbreak persistence in Equations (4)-(5). Here, is the number of beds in hospital and is the retail+recreation Google mobility metric for the county where hospital is located averaged across week (recall it is a daily metric). We lagged the mobility metrics by one month to account for delays between infection, hospitalization and reporting, as well as the time needed for the effects of a spike in mobility to trickle down into the general population. For counties with missing mobility metrics, we substituted the metrics from the nearest county with a similar urban makeup. The binary covariate was 1 if 3 months after the introduction of the Alpha variant in Quebec, introduced December 29, 2020, and the Omicron variant, introduced November 29, 2021, and 0 otherwise (NCCID, 2022). The introduction of the variants are marked in Figure 2(a) and we will in general refer to the outbreak that occurs in most hospitals around the introduction of Omicron as the Omicron outbreak. We included only in the relative odds of outbreak emergence and in the outbreak transmission rate to capture the short-term effects of the introduction of a new variant on outbreak emergence and transmission. We combined the Omicron and Alpha variants since there is likely not enough information in the data to estimate the effects of each separately. Additionally, having separate effects for the two variants can make it challenging to use the model in real-time, since we may not know if a future variant is more like Alpha or Omicron. Note, we did not have county-level information on the introduction of the new variants so is the same for all hospitals.
From a sample of patients from hospitals and grouped into neighborhoods, which divide Quebec, we calculated the spatial weight , in Equations (4)-(5), as
where, for example, is the proportion of sampled patients from hospital who lived in neighborhood , which is known as the Bhattacharyya coefficient (Bi et al., 2019). The Bhattacharyya coefficient must be between 0 and 1 and is a way to measure the amount of overlap between two categorical samples. Therefore, hospitals with more overlap in their catchment areas were given a larger weight. We took the 5 nearest neighbors of hospital , in terms of the largest weights, to form the neighborhood set . Note can be seen as a distance-based decay weight as where is the Bhattacharyya distance (Bi et al., 2019). To finish model specification, we used a uniform initial state distribution in each area.
We fit the CMSNB(1,2,4) model specified above to the Quebec hospitalization data using our proposed hybrid Gibbs sampler from Section 3. We ran the Gibbs sampler for 200,000 iterations on three chains with an initial burn-in of 50,000 iterations. All sampling was started from random values in the parameter space to avoid convergence to local modes. Convergence was checked using the Gelman–Rubin statistic (all estimated parameters<1.05), the minimum effective sample size (>1000) and by visually examining the traceplots (Plummer et al., 2006).
For comparison purposes we also fit a model without neighboring outbreak indicators in the transition probabilities, that is, with for in (4)-(5), which we will refer to as the Non-coupled Model. Additionally, we fit a two-state model without any absence or clone states, which we will refer to as the No Absence/Clone State Model. We chose these two models for comparison to examine the effects of our main contributions. We will sometimes refer to our CMSNB(1,2,4) model as the Full Coupled Model during comparison. Finally, we also compared to an endemic-epidemic (EE) model (Bracher and Held, 2022), which lies outside our Markov switching framework. The EE model is a state-of-the-art multivariate autoregressive count time series model which is popular for modeling spatio-temporal infectious disease counts (Bauer and Wakefield, 2018; Ssentongo et al., 2021) and is commonly used as a benchmark (Bauer et al., 2016; Stojanović et al., 2019). The EE model is usually used for forecasting and for investigating associations between covariates and disease transmission (Ssentongo et al., 2021). We compared to the latest iteration of the EE model (Bracher and Held, 2022) which allows for multiple temporal lags, whose order we chose using the WAIC. More details, including the parameter estimates, are given in SM Section 4.
Table 1 shows the WAIC for the 4 considered models. The Full Coupled Model has the lowest WAIC which lends support to the inclusion of neighboring outbreak indicators in the transition probabilities and the addition of the absence/clone states. The Markov switching models all have a lower WAIC compared to the Endemic-epidemic model, however, we would argue the main advantage of the Markov switching models is that they offer more interesting and richer interpretations. For instance, the Endemic-epidemic model concludes that mobility had a strong positive association with overall disease transmission, see SM Table 3. From Section 5.2 below, the CMSNB(1,2,4) model broadly agrees with this conclusion but provides a deeper understanding of the effect; mobility mainly affects disease transmission during the outbreak and not endemic periods and mobility increases the risk of outbreaks emerging and persisting and reduces the chances of the disease going extinct in smaller areas. These richer covariate interpretations could help policymakers better identify useful interventions. Additionally, the Markov switching models can be utilized for retrospective and real-time state estimation/forecasting, as we illustrate in Sections 5.3 and 5.4.
| Model | WAIC |
| Full Coupled | 17,516 |
| Non-coupled | 17,639 |
| No Absence/Clone | 17,545 |
| Endemic-epidemic | 17,845 |
5.2 Results
| Probability or (relative) Odds Ratio | ||||
| Disease | Disease | Outbreak | Outbreak | |
| Covariate | Emergence | Extinction | Emergence | Persistence |
| Intercept | .33 | .03 | .02 | .88 |
| (.15, .61) | (.01, .06) | (.01, .03) | (.83, .93) | |
| beds (100s) | 1.71 | .44 | 1.15 | 1.13 |
| (.82, 3.44) | (.21, .74) | (.86, 1.49) | (.85, 1.48) | |
| mobility ( 20%) | .77 | .53 | 1.73 | 1.65 |
| (.38, 1.36) | (.24, .98) | (1.08, 2.69) | (1.18, 2.26) | |
| new variant | – | – | 16.22 | – |
| (6.91, 33.14) | ||||
| weighted neighborhood | 1.15 | 1.01 | 1.98 | 1.31 |
| outbreak sum (.56) | (.84, 1.53) | (.73, 1.35) | (1.58, 2.46) | (1.14, 1.51) |
Table 2 gives the estimated parameters from the Markov chain part of the fitted CMSNB(1,2,4) model. As relative odds ratios from multinomial logistic regression can be difficult to interpret, we also plot some of the transition probabilities versus some of the covariates in Figure 4. Smaller areas were associated with a higher probability of COVID-19 extinction, which follows well known theories from Bartlett (1957) that state population size is inversely related to the rate of extinction of an infectious disease. We did not find evidence that larger areas experience more frequent or longer outbreaks compared to smaller areas (although outbreaks in larger areas do tend to be more severe in terms of transmission, see Table 3). We found that mobility in retail and recreation venues had a positive association with both outbreak emergence and outbreak persistence and a negative association with disease extinction. For example, we estimated that a 20 percent (one standard deviation) increase in the number of visits to retail and recreation venues, from the baseline week, was associated with a 73 (8, 169) percent increase in the odds of an outbreak emerging relative to remaining in the endemic state. However, this result may not imply a causal relationship between mobility and outbreak emergence. Mobility could be high at the start of an outbreak because outbreaks typically begin after a long period of relative calm, so it is natural for people to start going out more to movie theaters, restaurants, etc. It is difficult to control for this in our framework, as we would have to account for the time since the last outbreak ended, making the model non-Markovian. Also, note the high amount of uncertainty around the odds ratios for mobility.
The introduction of the Alpha and Omicron variants had a very strong association with the emergence of new COVID-19 outbreaks. We estimated that the probability of an outbreak emerging, at average levels of the other covariates and assuming no neighboring outbreaks, was .02 (.01, .03) if no new variant had been introduced recently and .19 (.09, .32) if a new variant had been introduced in the last 3 months. Emerging variants can trigger new COVID-19 waves due to them potentially being more contagious and more resistant to existing vaccines compared to previous variants (Maslo et al., 2022). Finally, we found that the presence of outbreaks in neighboring areas was strongly associated with outbreak emergence and persistence. We estimated that an outbreak occurring in a single neighboring area of average connectivity was associated with a 98 (58, 146) percent increase in the odds of an outbreak emerging relative to remaining in the endemic state and a 31 (14, 51) percent increase in the odds of an outbreak persisting. This could be due to either direct disease spread from the neighboring area, the signaling of spread from a common source and, potentially, partly due to correlated missing covariates. Interestingly, we found no strong evidence of an association between neighboring outbreaks and disease emergence or disease extinction, despite there being strong epidemiological justification for a relationship due to outbreak spread (Grenfell et al., 2001). Recall that the weighted sum of neighboring outbreak occurrence, in (4)-(5), is a latent covariate and so there is likely not a lot of information about its effect on these rarer transitions (the disease only goes extinct in the smaller catchment areas).
| Rate Ratios | |||
| Covariate | Parameter | Endemic | Outbreak |
| Intercept of random intercepts | .99 | 2.18 | |
| (.85, 1.13) | (1.97, 2.41) | ||
| Std. dev of random intercepts | .31 | .13 | |
| (.21, .44) | (.1, .17) | ||
| beds (100s) | 1.07 | 1.20 | |
| (1.02, 1.11) | (1.09, 1.32) | ||
| mobility (20%) | 1.06 | 1.17 | |
| (1, 1.13) | (1.14, 1.20) | ||
| new variant | – | 1.01 | |
| (.97, 1.06) | |||
| autoregressive | .69 | .75 | |
| (.65, .72) | (.72, .78) | ||
| overdispersion | 5.50 | 8.95 | |
| (4.13, 7.37) | (7.89, 10.14) | ||
Table 3 gives the estimated parameters from the count part of the model. Transmission during the outbreak periods was on average 121 (91, 157) percent higher than during the endemic periods. Mobility in retail and recreation venues and population size were both positively associated with transmission during the outbreak and endemic periods, and outbreaks amplified the effects of population size and mobility on transmission. Note, as we are modeling hospitalizations, an effect in terms of transmission in the hospitalizations could reflect an effect on transmission in the actual cases and/or an effect on the severity of the disease. Mobility and beds should not affect disease severity; however, new variants often do (Chenchula et al., 2022). This could partly explain why we found the introduction of a new variant likely did not have a large effect on transmission in the hospitalizations, Omicron was much more transmissible person to person but less severe (NCCID, 2022). Also, we combined the Alpha and Omicron variants, and, as can be seen in Figure 2, Alpha does not appear to have had a large impact on transmission in the hospitalizations. Finally, for sensitivity analysis, we also fit the CMSNB(1,2,4) model using .05 and .1 in place of .01 in the constraint (6). The only posterior that changed noticeably was for the effect of mobility on transmission in the endemic period () which increased from 1.06 (1, 1.13) to 1.08 (1.02, 1.15) going from .01 to .1 in the constraint.
5.3 Retrospective evaluation and comparison
As mentioned in Section 3.1 it is important to examine the retrospective state estimates of an epidemiological Markov switching model, that is, the posterior probability that the disease was in each state during each week of the study period, to ensure the models’ estimates of the epidemiological history of the disease are sensible. Starting with the bottom graphs of Figures 5 (a) and (c), we compare the retrospective state estimates for Fleury Hospital in Montreal, one of the smaller hospitals, between models with, (a), and without, (c), an absence state, as it is a good example of how accounting for long periods of disease absence in smaller areas can be important. The top graphs show the posteriors of the endemic and outbreak state distributions, which can be helpful for better understanding the state estimates. To draw from the posterior of the state distributions we drew from , see (1)-(2), for , , (endemic) and (outbreak). The Full Coupled Model identifies 3, unlikely 4, outbreak periods for Fleury Hospital which are separated by a few endemic and absence periods. Long sequences of zeroes in Fleury Hospital are generally assigned to absence periods by the Full Coupled Model. In contrast, the No Absence/Clone State Model assigns the long strings of zeroes to endemic periods, bringing the endemic state distribution much closer to 0 compared to the Full Coupled Model. This causes some issues in the state estimation for the No Absence/Clone State Model, as the outbreak state becomes too dominant. Firstly, the No Absence/Clone State Model classifies a very small increase in hospitalizations, around week 56, as likely a short outbreak period, which is not realistic. Secondly, the No Absence/Clone State Model identifies a smaller outbreak, around week 81, before the final Omicron outbreak. In contrast, the Full Coupled Model only identifies the Omicron outbreak here, which appears more realistic as the number of hospitalizations in the few weeks before Omicron never becomes high enough to be very concerning and, by checking the other plots, there is no strong evidence of outbreaks in any neighboring areas.
In Figures 5 (b) and (d) we compare the retrospective states estimates for Hôpital de Verdun between models with, (b), and without, (d), clone states. The No Absence/Clone State Model shows rapid switching between the outbreak and endemic states around weeks 50 and 102. During week 50, it is more realistic that the one-week dip in the hospitalizations is due to random variation, which is what the Full Coupled Model identifies, and that there are not multiple outbreaks occurring. The Full Coupled Model does place some weight on an endemic period between weeks 102 and 103. A two-week dip into the endemic period seems more realistic than a one-week dip, however, we could prevent two-week dips by increasing the number of endemic clone states. For the Full Coupled Model, we looked through the plots in Figure 5 for each hospital and found the retrospective state estimates to be sensible.
5.4 Real-time evaluation and comparison
To evaluate the real-time performance of the models we fit the Full Coupled and Non-coupled models up to 6 weeks before the introduction of Omicron and then up to every week after, that is, we fit the models up to time for , producing 30 sets of posterior samples for each model. We did not include as a covariate since many of the fitted models are fit before, or just after, Omicron and so there is likely not enough information to estimate the effect for many of the fits. Additionally, we found after discarding 27% of the data for the real-time evaluation that MCMC chains would sometimes get stuck in local modes in regions of the parameter space that only used the absence state and one of the count states. This is not an uncommon problem in complex Bayesian mixture modeling and, following the discussion and recommendations in Section 4.2.3 of Frühwirth-Schnatter (2006), we applied mild shrinkage to the transition probabilities by assigning priors to the intercepts in Equations (4)-(5), which stabilized model fitting. Figure 6 compares the posterior probabilities that an outbreak is currently happening (bottom graphs solid lines) between the Full Coupled Model, in orange, and the Non-coupled Model, in purple, for four hospitals. Recall from Section 3.1 that the posterior probability that an outbreak is currently happening is given by and would be used for the purpose of outbreak detection in a real-time scenario. The Full Coupled Model gives an earlier warning of the Omicron outbreak in the first three hospitals, (a)-(c), but not in the final hospital, (d). In (d) the hospitalizations rise very rapidly and so there is a lot of within area information about the Omicron outbreak, making the coupling less useful for providing an early warning. SM Figure 5 (a) shows that, when averaged across all hospitals, early on during the Omicron wave the Full Coupled Model gives a .15-.2 higher posterior probability that an outbreak is currently happening, an earlier warning compared to the Non-coupled Model, while it does not show an increased risk of an outbreak prior to Omicron. The Omicron wave was highly synchronized between the hospitals and, therefore, as can be seen from SM Figure 5 (b), the probability of outbreak emergence would have been higher for the Full Coupled Model during the start of the wave as evidence of outbreaks accumulated in neighboring hospitals. In general, SM Figure 5 shows that the Full Coupled Model tends to enforce the status quo, making outbreak occurrence more (less) likely if there are (are not) outbreaks occurring in neighboring areas.
The dashed line in the bottom graphs of Figure 6 show the one-week ahead outbreak forecasts , from the previous week, for the Full Coupled Model. Note that, we typically observe around the start of the Omicron wave. We looked through the plots in Figure 6 for each hospital and found the real-time state estimates/forecasts to be sensible in most hospitals. We did find evidence of false alarms being thrown in 2/30 hospitals during the real-time evaluation, shown in the SM Section 7.3. The false alarms appear to have been mostly caused by hospitalizations spiking during periods of high outbreak risk, meaning the model has no information that an outbreak starting should be unlikely to counter the false alarm. Still, the model corrected itself quickly, by the next week in most cases.
Finally, the map in SM Figure 6 shows the likely state in the catchment area of each hospital during the last week of the study period, March 5th, 2022, from the Full Coupled Model with new variant as a covariate. According to the map, most catchment areas were in the outbreak state at this time, especially in/around Montreal.
6 Discussion
We have proposed a three-state coupled nonhomogeneous Markov switching model for the general analysis of spatio-temporal outbreak occurrence. The model can be used for investigating associations between various factors, including geographical outbreak spread, and outbreak emergence, outbreak persistence, disease extinction and disease emergence, as well as for detecting and forecasting the outbreaks. We made three main contributions to the existing (Amorós Salvador, 2017) endemic/outbreak Markov switching literature. Firstly, to account for long periods of disease absence in smaller areas, (Bartlett, 1957) we added an absence state to our model in addition to the more traditional endemic and outbreak states. Secondly, to incorporate geographical outbreak spread (Grenfell et al., 2001), we allowed the transition probabilities to depend on whether outbreaks were occurring in neighboring areas. Previous two-state approaches (Heaton et al., 2012) have allowed the probabilities of outbreak emergence to depend on neighboring outbreaks, but they did not allow outbreak emergence to depend on covariates, and they fixed the probability of outbreak persistence at one to prevent rapid switching between endemic and outbreak periods, meaning their method could not be applied to multiple outbreaks. Finally, to allow for the analysis of multiple outbreaks, we instead introduced clone states (Kaufmann, 2018) into the model, which prevents rapid switching between endemic and outbreak periods by enforcing a minimum endemic and outbreak duration.
We applied our model to the analysis of COVID-19 outbreaks across Quebec based on admissions in the 30 largest hospitals. We found that mobility in retail and recreation venues, the introduction of new variants and the occurrence of outbreaks in neighboring areas all had positive associations with the emergence or persistence of new COVID-19 outbreaks in Quebec. Additionally, mobility and population size had positive associations with transmission during the outbreak and endemic periods, and the effects were amplified during the outbreak periods. As for disease extinction and emergence, we only found evidence that the disease is more likely to go extinct in smaller areas and at lower levels of mobility. Disease extinction and emergence are rarer transitions, so more data needs to be collected to study them with high precision.
Regarding model performance, we found our model gave realistic estimates of past epidemiological states for the Quebec hospitalizations and appeared to perform outbreak detection/forecasting well in a real-time evaluation of the Omicron wave, absent a small number of false alarms during the real-time evaluation. However, on the Quebec hospitalizations it was difficult to exactly quantify the accuracy of the state estimates, e.g., calculate AUC, sensitivity, etc., as the true underlying states were not known. Therefore, we also conducted a simulation study with known outbreaks and found our model achieved good sensitivity, specificity and timeliness both retrospectively and during real-time outbreak detection and forecasting.
Regarding model comparison among different Markov switching models. In our simulation study, where we assumed outbreaks in neighboring areas occurred around the same time, including neighboring outbreak indicators in the transition probabilities greatly improved sensitivity, specificity and timeliness during real-time outbreak detection and forecasting. As for our analysis of the Quebec hospitalizations, model comparison using WAIC supported the inclusion of absence/clone states and the addition of neighboring outbreak indicators in the transition probabilities. We also found for the Quebec analysis that the addition of an absence state can be important in smaller areas for reducing bias in the endemic state distribution towards 0, a bias that we observed often leads to unrealistic state estimates. We further observed that clone states are important for preventing rapid switching between endemic and outbreak periods, making the state estimates more realistic. Finally, in a real-time evaluation we found that the incorporation of outbreak spread led to an earlier warning of the Omicron wave in Quebec as the wave was highly synchronized across the hospitals. Ultimately, our contributions show a lot of promise for improving state estimation both retrospectively and in real-time, especially when there are small areas and highly spatially synchronized outbreaks. Although we applied our model to spatio-temporal counts of hospitalizations, it could also be applied to spatio-temporal counts of disease cases, which are popular to model for outbreak occurrence analysis (Knorr-Held and Richardson, 2003; Watkins et al., 2009), or deaths.
Now we will address the question of how our approach fits into the broader popular literature on spatio-temporal infectious disease modeling, including compartmental models (Bauer and Wakefield, 2018) and multivariate time series models (Ssentongo et al., 2021). One aspect of our approach that we believe is unique is the ability to quantify associations, through odds ratios, between covariates and certain epidemiological transitions, such as outbreak emergence or disease extinction. Also, our model can provide the probability that a spike in an epidemiological indicator is due to an actual outbreak developing as opposed to just random endemic variation, which is valuable for outbreak detection (Unkel et al., 2012). On the Quebec hospitalizations, we found our model fit better and had better interpretations compared to the Endemic-epidemic model (Bracher and Held, 2022), a popular alternative.
There are also some limitations with our approach. Conditional on the states, we only consider first-order autoregression, see the discussion at the end of SM Section 4. Given weekly data and a disease with a short serial interval, such as COVID-19, this is likely not a major limitation (Bracher and Held, 2022). However, in some cases, especially with daily data, it can be important to account for higher-order temporal dependencies. As another limitation, we assume that the probability of an outbreak emerging does not depend on the amount of time since the last outbreak has ended. After an outbreak ends, the susceptible population should increase over time due to demographic changes, waning immunity and other factors, leading to an increased outbreak risk (Keeling and Rohani, 2007). In a hidden semi-Markov model (Langrock and Zucchini, 2011), the transition probabilities can depend on the amount of time that has been spent in the state, which can be pursued in further work. Finally, given the large number of transitions in our model, the number of covariate effects grows quickly with the number of covariates (4 times). Shrinkage could be applied to the covariate effects in the transition probabilities to eliminate many unimportant effects (Wang et al., 2023).
Acknowledgements
This research was part of a larger project of INESSS (Institut national d’excellence en santé et en services sociaux), whose objective was to produce bed occupancy projections for COVID-19 patients. The access to data was made possible through a tripartite agreement between the MSSS, the RAMQ and INESSS. This work is part of the PhD thesis of D. Douwes-Schultz under the supervision of A. M. Schmidt in the Graduate Program of Biostatistics at McGill University, Canada. Douwes-Schultz is grateful for financial support from IVADO and the Canada First Research Excellence Fund/Apogée (PhD Excellence Scholarship 2021-9070375349). Schmidt is grateful for financial support from the Natural Sciences and Engineering Research Council (NSERC) of Canada (Discovery Grant RGPIN-2017-04999). Shen is grateful for financial support from the Fonds de recherché du Québec–Santé (FRQS) (Doctoral training award 313602). Buckeridge is supported by a Canada Research Chair in Health Informatics and Data Science (950-232679. This research was enabled in part by support provided by Calcul Québec (www.calculquebec.ca) and Compute Canada (www.computecanada.ca).

### v3

@@
 For example, dengue outbreaks are unlikely to occur during the winter when mosquito activity is low (Descloux et al., 2012), the introduction of new variants, such as Omicron, has likely played an important role in the development of new COVID-19 waves (Maslo et al., 2022) and disease extinction is less likely in smaller communities (Bartlett, 1957).
+Note it can often make sense to include the same covariate in both and .
+For instance, as mentioned, population size should be related to disease extinction, however, it may also affect disease transmission during the endemic and outbreak periods as individuals are more likely to move through high-population areas (Grenfell et al., 2001).
 As for the inclusion of neighboring outbreak indicators in (4)-(5), individuals will mix with those in other areas which can cause outbreaks to spread geographically, a phenomenon known as traveling waves (Grenfell et al., 2001).
@@
 5.4 Real-time evaluation and comparison
-To evaluate the real-time performance of the models we fit the Full Coupled and Non-coupled models up to 6 weeks before the introduction of Omicron and then up to every week after, that is, we fit the models up to time for , producing 30 sets of posterior samples for each model.
+To evaluate the real-time performance of the models we fit the Full Coupled, Non-coupled and EE models up to 6 weeks before the introduction of Omicron and then up to every week after, that is, we fit the models up to time for , producing 30 sets of posterior samples for each model.
 We did not include as a covariate since many of the fitted models are fit before, or just after, Omicron and so there is likely not enough information to estimate the effect for many of the fits.
@@
 This is not an uncommon problem in complex Bayesian mixture modeling and, following the discussion and recommendations in Section 4.2.3 of Frühwirth-Schnatter (2006), we applied mild shrinkage to the transition probabilities by assigning priors to the intercepts in Equations (4)-(5), which stabilized model fitting.
-Figure 6 compares the posterior probabilities that an outbreak is currently happening (bottom graphs solid lines) between the Full Coupled Model, in orange, and the Non-coupled Model, in purple, for four hospitals.
+Figure 6 compares the posterior probabilities that an outbreak is currently happening (bottom graphs solid lines) between the Full Coupled Model, in red, and the Non-coupled Model, in blue, for four hospitals.
 Recall from Section 3.1 that the posterior probability that an outbreak is currently happening is given by and would be used for the purpose of outbreak detection in a real-time scenario.
@@
 Still, the model corrected itself quickly, by the next week in most cases.
+The top graphs of Figure 6 summarize the one-week ahead predictions, from the previous week, of the hospitalizations from the Full Coupled and Non-coupled models.
+SM Figure 9 summarizes the one-week ahead predictions from the EE model for the same hospitals.
+In panels (a)-(c) of Figure 6, there is a much larger discrepancy between the state estimates of the two models than between the predictions of the counts.
+This indicates that differences in predictive performance may be a poor proxy for differences in the accuracy or, especially, timeliness of the state estimates.
+Outbreaks often emerge at low counts where there is a lot of overlap between the state distributions, see Figure 5.
+Therefore, a large difference in the state estimates during the start of an outbreak may not translate to a large difference in predictions.
+We used the multivariate log score (Bracher and Held, 2022), see SM Section 7.4, to compare the one-week ahead forecasts from the three models.
+The Full Coupled Model obtained the lowest mean score, indicating the best performance on average across the entire Omicron outbreak, at 3.20.
+This was followed by the Non-coupled, 3.25, and EE, 3.27, models.
+Permutation tests (Paul and Held, 2011) showed moderate evidence of a true difference in the overall mean scores, with two-sided p-values of .033, Full Coupled versus Non-coupled, and .052, Full Coupled versus EE.
+To investigate local differences in forecasting performance, we plot the multivariate log scores across time in SM Figure 10.
+The Full Coupled Model obtained a much lower score, indicating superior predictive performance, around the emergence of the Omicron wave, which is the most important period to capture for disease surveillance (Buckeridge, 2007).
+There are some important limitations to using the multivariate log scores to compare the models.
+We can only evaluate the models over the Omicron wave, where model fitting is stable, which includes just 29 weeks and is missing some important features of the full data set such as long strings of zeroes.
+In contrast, the WAIC evaluates the models over the entire data set.
+Finally, an important advantage of the Markov switching models for forecasting, over the EE model, is that they can determine whether a forecasted increase in hospitalizations will likely be due to an outbreak emerging, as opposed to random endemic variation.
+See the bottom graphs in Figure 6.
+If an increase in the counts is due to endemic variation it will be much less concerning to health authorities (Unkel et al., 2012).
 Finally, the map in SM Figure 6 shows the likely state in the catchment area of each hospital during the last week of the study period, March 5th, 2022, from the Full Coupled Model with new variant as a covariate.
@@
 As another limitation, we assume that the probability of an outbreak emerging does not depend on the amount of time since the last outbreak has ended.
-After an outbreak ends, the susceptible population should increase over time due to demographic changes, waning immunity and other factors, leading to an increased outbreak risk (Keeling and Rohani, 2007).
+After an outbreak ends, the susceptible population should increase over time due to demographic changes, waning immunity, and other factors, leading to an increased outbreak risk (Keeling and Rohani, 2007).
 In a hidden semi-Markov model (Langrock and Zucchini, 2011), the transition probabilities can depend on the amount of time that has been spent in the state, which can be pursued in further work.
-Finally, given the large number of transitions in our model, the number of covariate effects grows quickly with the number of covariates (4 times).
+Additionally, given the large number of transitions in our model, the number of covariate effects grows quickly with the number of covariates (4 times).
 Shrinkage could be applied to the covariate effects in the transition probabilities to eliminate many unimportant effects (Wang et al., 2023).
+Finally, we did not include vaccination data in the model.
+While data on the percentage of individuals with at least two doses is available in Quebec (INSPQ, 2022), there is very little variation in the covariate.
+Vaccination coverage increased rapidly and was only greater than zero during the final Omicron outbreak.
+Also, there are likely interactions between vaccination and new variants (Chenchula et al., 2022).
+Therefore, we decided there was not enough information in the available data to capture the complicated effects of vaccination with the model.
+Not including vaccination information could have consequences on the estimated effects of mobility and new variants.
+For instance, there were fewer mobility restrictions on vaccinated individuals (Cabrera and Olson, 2021) who were also less likely to contract, especially severe, cases of the disease (Johnson, 2022).
 Acknowledgements